a boat can travel 5 miles downstream in the same length of time it travels 3 miles upstream. if the rate of the current is 3mph, find the speed of the boat in still water.
d=5 ?
r=3mh ?
t=?
so simple yet a little hard......????

Answer by stanbon(26269)

a boat can travel 5 miles downstream in the same length of time it travels 3 miles upstream. if the rate of the current is 3mph, find the speed of the boat in still water.
-------
Downstream DATA:
distance = 5 miles ; rate = b+3 mph ; time = d/r = 5/(b+3) hrs
----------------
Upstream DATA:
distance = 3 miles ; rate = b-3 mph ; time = d/r = 3/(b-3)
=============================================================
Equations:
time up = time down
3/(b-3) = 5/(b+3)
---
3(b+3) = 5(b-3)
3b+9 = 5b-15
2b = 24
b = 12 mph (speed of the boat in still water)
===============================================
Cheers,
Stan H.